nh2coonh4---- 2nh3 co2 in a 2 litre vessel.initially 5 moles of nh2coonh4 are present and it dissociates 40% to reach equilibrium . 1 mole of co2 is removed .calculate new concentrations

Grade:11

1 Answers

Ujjawal
5 years ago
NH2COONH4--------------------------------2NH3+CO2
Conc. =n/v=5/2                                    0          0
Final concert=5/2-40/100 ×5/2         2           1
                           =5/2-1
                              3/2
Final CO2 conc is 1M mean 1mole in 1litre but volume is 2 litre so m1v1=m2v2 so m2comes to be 1/2 molar which is final concert of CO2 .
 
After removing 1mole CO2 no CO